# Predicting Census Survey Response Rates with Parsimonious Additive Models and Structured Interactions

### v4

@@ 3 Incorporating strong hierarchy constraints
In this section, we discuss the model with sparse interactions under strong hierarchy.
@@
For example, based on our analysis, covariates “Males” and “Females” appear to be unnecessary for self-response prediction when the strong hierarchy constraint is imposed via model (3), while a related covariate “Households with female without spouse” remains important. In contrast, the LRS uses both “Males” and “Households with female without spouse” (Erdman and Bates, 2016) – both features appear to be important perhaps due to the use of a linear regression model. Similarly, covariates pertaining to civilian employment status across different segments/age-groups of the population, are excluded from our nonlinear AM when we learn the (sparse) main and interaction effects under strong hierarchy.
@@
To our knowledge, ours is the first work to study estimator (3). The methodology presented here is of independent interest in the context of structured nonparametric learning with interactions. We refer to this model as ELAAN-H [^12] ELAAN-H stands for End-to-end Learning Approach for Additive spliNes with Hierarchy.. In Section S3 of the Supplement, we propose algorithms to obtain good solutions to Problem (3).
@@
As mentioned earlier, the EBM approach (Nori et al., 2019) does not support hierarchy constraints. GAMI-Net (Yang, Zhang and Sudjianto, 2021) follows a two-stage approach to fit nonparametric additive models with sparse (weak) hierarchical interactions. GAMI-Net performs a screening step after estimating the main effects with neural-network blocks and only consider interaction effects that satisfy the weak hierarchy principle amongst the screened main effect – an interaction effect can appear in the model if one of the main effects is in the model. As mentioned earlier, GAMI-Net is not based on an optimization procedure; and can be computationally much more expensive than our estimators.
@@ 4.1 Sparse additive model with interactions
We measure prediction accuracy using the integrated squared error, $\text{ISE}=\mathbb{E}_{\mathbf{x}}[(\hat{f}(\mathbf{x})-f^{*}(\mathbf{x}))^{2}]$, estimated by Monte Carlo integration using $10,000$ test observations from the same distribution as the training ones, following the procedure in Li, Lue and Chen (2000). We vary the number of training observations from $100$ to $400$ and use $100$ replications for each simulation setting. We compare our estimators to well-known benchmarks EBM (Nori et al., 2019), GAMI-Net (Yang, Zhang and Sudjianto, 2021), COSSO (Lin and Zhang, 2006), and MARS [^13] MARS is a stepwise forward?backward procedure for building functional ANOVA models. (Friedman, 1991).
@@
Table 2 presents the average test ISE and their standard errors (based on the replications). The tuning parameters for ELAAN-I, ELAAN-H, EBM and GAMI-Net are selected via 5-fold cross-validation. The results for MARS and COSSO are taken from Table 6 in Li, Lue and Chen (2000). Table 2 demonstrates that both of our estimators, ELAAN-I and ELAAN-H, outperform the competitors across the entire range of training set sizes under consideration. Due to space constraints, additional simulations comparing our approach to EBM and with similar conclusions are provided in the Supplement Section S4.2.2.
Next, we study the support recovery performances of ELAAN-I, ELAAN-H, EBM and GAMI-Net. We evaluate models using discrepancy between the true and estimated support and separately report the average F1-score for the main and the interaction effects. The support recovery metrics are shown in Table 2. We observe that, overall, both ELAAN-I and ELAAN-H appear to outperform EBM and GAMI-Net in terms of the F1-score. Moreover, ELAAN-H is seen to be the best-performing method across all the simulation settings. In summary, our approaches work quite well in terms of variable selection.
@@
Table 3 presents additional variable selection details of the two best-performing approaches: ELAAN-I and ELAAN-H. For each true main and interaction effect, we report the relative frequency of their appearance in the estimated support across all replications. The results in Tables 2 and 3
suggest that ELAAN-H outperforms ELAAN-I in terms of the variable selection when the true model is hierarchical. In practice, when we do not know if the underlying truth obeys strong hierarchy, ELAAN-H tends to result in more compact models, as we demonstrate in the Census application in Section 5.
@@ 4.2 Large-scale setting with correlated features
ELAAN-H ${}~{}10.6\pm 1.8$ ${}~{}97.57\pm 0.68$ $\textbf{97.57}\pm 0.68$ $91.31\pm 1.52$
We produce 10,000 training observations and evaluate the prediction performance as before, using ISE on a test set of size 10,000. We compare our estimators to EBM (Nori et al., 2019) and GAMI-Net (Yang, Zhang and Sudjianto, 2021). Table 4 presents the average prediction errors and support recovery metrics over 25 simulation replications, along with the corresponding standard errors. The tuning procedure for ELAAN-I, ELAAN-H, EBM and GAMI-Net is described in the Supplement. Table 4 demonstrates that our proposed models consistently and significantly outperform the competitors, exhibiting a 7-11 fold reduction in the prediction error.
@@ 5 Case study: Predicting the Census Survey Self-Response Rate
We now present our findings on the Census application. As mentioned in Section 1, obtaining interpretable models with good predictive capabilities is important in this application – such models can inform the planning of outreach campaigns and guide stakeholders in deciding the allocation of spending for different communications channels (for example, TV, radio, digital), advertising messages with appropriately tailored content, and the timing of spending during the campaigns (Kulzick et al., 2019). Due to their opaque nature, predictive models that topped the nationwide competition were not actionable for this application. On the other hand, the linear models that drive the LRS have limited predictive power. We explore how our proposed approach balances simplicity and good prediction performance.
The data we use for this study is publicly available in the US Census Planning Database, which provides a range of demographic, socioeconomic, housing, and Census operational data (US Census Bureau, 2019). The data in the Planning Database includes covariates from the 2010 Census and the 2013-2017 American Community Survey (ACS), aggregated at both the Census tract level and the block level. We use tract-level data, with approximately $74,000$ observations and $500$ covariates. The response is the ACS self-response rate. We exclude the following covariates from our model: spatial covariates (“State”, “County”, “Tract”, “Flag”, “AIAN Land”) and variables that serve as a proxy to the response (for example, “Low response score”, “Number of housing units that returned first forms”, “Replacement forms” or “Bilingual forms” in Census 2010). We also remove the margin of error variables corresponding to the ACS. After excluding these variables, we are left with $p=295$ covariates [^14] To clarify, the results presented in this section are based on these $295$ features, though our approach scales to $p=500$ features..
@@ 5.1 Experimental setup
Benchmark Methods. We evaluate and compare our models against existing linear and nonparametric approaches. We consider the following linear models with main effects only (i.e., without interactions): (i) Ridge regression; (ii) Lasso regression (Hastie, Tibshirani and Wainwright, 2015); (iii) $\ell_{0}$-penalized regression with additional ridge regularization (denoted as $\ell_{0}-\ell_{2}$), implemented via L0Learn (Hazimeh and Mazumder, 2020a). In addition, we consider the following linear models with interaction effects: (iv) Lasso with main effects and all pairwise linear interactions, (v) hierScale: a convex optimization framework for learning sparse main-effects and interactions under a strong hierarchy constraint (Hazimeh and Mazumder, 2020b). For (i), (ii), and (iv), we use Python’s scikit-learn library (Pedregosa et al., 2011).
In addition to the competing methods discussed above based on linear models, we consider nonparametric additive models with/without interactions. For AMs with main effects only, we compare with (vi) Additive Models with Lasso-based selection, which we fit using Python’s scikit-learn library (Pedregosa et al., 2011). For additive models with interactions, we consider (vii) EBM: A boosting approach that uses trees for main and interaction effects (Nori et al., 2019). (viii) GAMI-Net: a neural network-based additive model with structured interactions (Yang, Zhang and Sudjianto, 2021). For EBM, we tune the learning rate in the range $[10^{-4},10^{-1}]$ and the number of interactions in the range $[10,500]$ for 1000 trials. For GAMI-Net, we run the model once with a $200$ batch-size, $0.0001$ learning rate, $500$ interaction effects, $500$ epochs (for each stage) and $0.0001$ loss threshold. We note that GAMI-Net took 3 days to complete using a machine wth 8-CPUs and 128GB RAM.
@@
All our algorithms for estimators (a)–(c) are implemented in Python. We use cubic B-splines with $10$ knots for the main effects. For the interaction effects, we use tensor spline bases of degree $3$ with 5 knots in each coordinate, leading to a total of $5\times 5=25$ knots [^15] We study the effect of number of knots on out-of-sample generalization on synthetic data (see Supplement Section S4.2.1).. Because the problem at hand has $295$ covariates and $43,365$ possible pairwise interactions, we need to be careful with implementation aspects while generating spline-transformed representations for all the interaction effects, which can be memory intensive.
For ELAAN and ELAAN-I we perform a tuning procedure with warm-starts over a 2D grid of parameters ($\lambda_{1}$, $\lambda_{2}$) – for details, see Supplement Section S1.4.3. For both ELAAN-I and ELAAN-H, we set $\alpha=1$ so that the main and the interaction effects have the same $\ell_{0}$-penalty parameter. For ELAAN-H, we fix the smoothing parameter $\lambda_{1}$ to the optimal value available from ELAAN-I. Parameter $\lambda_{2}$ (as well as parameter $\tau$ defined in Supplement Section S3) is chosen based on validation tuning. For all our methods (ELAAN, ELAAN-I, ELAAN-H), we cap the number of tuning-parameter values at 1000.
@@ 5.2 Comparing methods: prediction, sparsity and model structure
Table 5: Comparisons of our methods with several benchmark models as discussed in the text. We display the average test RMSE for the different models, along with the corresponding number of covariates, and number of effects (main and interactions). Best metrics are highlighted in bold. Numbers after $\pm$ provide the standard errors. Asterisk(*) indicates statistical significance (p-value¡0.05) over the best existing model, using a one-sided paired t-test. Models (ix), (x) are black-box models using higher order interactions, hence #Main and #Interactions are left as ‘-’ (dash).
@@
Additive models vs black-box ML methods. Table 5 reports the prediction errors (RMSE) on the test-set along with the number of nonzero features in the model. Importantly, we observe that the test performances of AMs with sparse interactions (both with and without the hierarchy constraints) are better than that of the best black box predictive ML models (ix), (x) that are based on GBDT, Neural Networks. Our model ELAAN-H delivers the best RMSE value and is closely followed by ELAAN-I.
ELAAN-I/ELAAN-H  vs prior art for AMs with sparse interactions.
Among the methods we compared, EBM and GAMI-Net were the only methods that could compute AMs with sparse interactions for the Census dataset. Interestingly, the test performances of both ELAAN-I and ELAAN-H are better than EBM and GAMI-Net. Interpretability (i.e., model sparsity and/or hierarchy) is another key factor differentiating the leading prediction methods. Our models select a smaller number of additive components than EBM and GAMI-Net. For example, ELAAN-I selects a total of 234 additive components. In contrast, EBM and GAMI-Net select 787 and 374 additive components, respectively. Similarly, we also observe that our models are substantially more compact in terms of the number of selected covariates by a factor of (approximately) two.
@@
We briefly discuss computation times for different methods. EBM takes around one hour to fit a model with 500 pairwise interactions (for a single tuning parameter). Computing a GAMI-Net solution (for one tuning parameter corresponding to 500 interaction effects) takes around $3$ days using 8 CPUs and 10 hours using a single V100 Tesla GPU. In contrast, our approaches are much faster: ELAAN-I takes on average $1.8$ mins to compute a solution for a fixed tuning parameter [^16] We use warm-start continuation across $\lambda_{2}$-values for a fixed $\lambda_{1}$: it takes approximately 3 hours to compute 100 solutions where the maximum number of pairwise interactions is 500.. When ELAAN-H is run on a reduced subset of 800800~{}800800 pairwise interactions, it takes $1.5$ hours to compute a path of 100 solutions. These numbers are reported on an 8-CPU 128GB RAM device.
@@
Nonlinear models vs linear models. Table 5 suggests that nonlinear/nonparametric models have better predictive performance than their linear counterparts. Also, linear models with sparse main and interaction-effects appear to have an edge over sparse linear models without interactions. Another appealing aspect of sparse nonparametric AMs compared to linear models, is in model parsimony – we alluded to this aspect in Figure 3 while using a reduced number of features. The number of nonzero effects is significantly lower for the nonlinear AMs. For example, $\ell_{0}$-sparse nonparametric AMs with no interactions (i.e., ELAAN) can achieve the same level of predictive performance as its linear model counterpart (i.e., Model (iii)) with significantly fewer covariates: $40$ versus $145$. Similarly, if we compare hierScale (i.e., linear main and interaction effects with strong hierarchical restrictions) to its nonparametric counterpart i.e., ELAAN-H, the number of main effects (and also, covariates) reduces from $276$ to $133$, and the number of interaction effects reduces from $4,107$ to $255$. Interestingly, all the covariates selected by the nonparametric AMs with interaction models (with/without strong hierarchy) are contained in the set of covariates selected by the sparse linear models with both main and interaction effects.
@@
Additive nonlinear models: interactions vs no interactions. Table 5 shows that ELAAN (AMs with no interactions) has more covariates than AMs with interactions, both ELAAN-I and ELAAN-H. By including interactions that obey the hierarchy principle, we select a smaller number of covariates: ELAAN-I has 154 covariates vs ELAAN-H has 133. The difference between ELAAN vs ELAAN-I is possibly because many nonlinear main effects attempt to explain the nonlinear interaction effects. This reduction points to the redundancy of some of the covariates when interaction effects are directly included in the model. The prediction performance of nonlinear AMs seems to improve significantly when pairwise interactions are included in the model.
@@
Hierarchy vs no hierarchy. We observe that ELAAN-H, i.e., additive model with hierarchical interactions, achieves the best out-of-sample RMSE – this improves over ELAAN-I (interactions with no hierarchy). The improvement is statistically significant based on a paired t-test. We observe that ELAAN-I selects fewer effects (33 Main + 201 Interactions) when compared to ELAAN-H (133 Main + 255 Interactions). On the other hand, ELAAN-H obeys the hierarchy constraint and selects fewer covariates than ELAAN-I– this can aid in interpretability and be easier to operationalize from a practical standpoint.
@@
Insights from visualizations. As illustrated by Figure 2, which displays some marginal nonlinear fits, an appealing aspect of nonlinear AMs is that they naturally allow the practitioner to gather insights into the associations between the response and a feature of interest, for example, by exploring the map $x_{j}\mapsto f_{j}(x_{j})$. Similarly, the map $(x_{j},x_{k})\mapsto f_{j,k}(x_{j},x_{k})$ would shed light into how the interaction of $(x_{j},x_{k})$ influences the output. Such interpretations are a by-product of our additive model framework; and may not be readily available via black-box ML methods such as Neural Networks and GBDT—see also Lou et al. (2013) for related discussions advocating for the interpretability of AMs. These association plots, for example, can help stakeholders identify promising factors and patterns influencing self-response scores potentially informing policy decisions (e.g, targeted investments and optimizing operational costs).
@@
To obtain a finer understanding of the performance of our models, we use visualization tools inspired by earlier works from the US Census Bureau (US Census Bureau, 2017; Kulzick et al., 2019). Figure 4 illustrates the tract self-response rates predicted by ELAAN-I i.e., model (12) on a map of the United States. Different from Figure 1, which shows actual data with missing values for some of the tracts [^17] Some responses are deliberately suppressed by the US Census Bureau for privacy considerations, to limit the disclosure of information about individual respondents and to reduce the number of estimates with unacceptable levels of statistical reliability (US Census Bureau, 2016), Figure 4 provides predicted self-response rates for all tracts based on our proposed model. Predictive models in general and our models in particular, allow us to forecast the response rates for tracts where gathered data is incomplete. It is worth noting that our regression-based models may be preferable over commonly used nearest neighbor type methods: the latter models do not generally offer insights into factors associated with high/low response rates. Nearest neighbor based methods may also be ill-posed for isolated regions such as Alaska, Hawaii and Puerto Rico, each having at least $10\%$ of their tracts censored.
@@
Figure 6 shows how the sorting of the Census tracts into quintiles (US Census Bureau, 2017; Kulzick et al., 2019) of the actual self-response rates compares with the corresponding sorting of the predictions made by: the Lasso (we use an $\ell_{1}$-penalized linear model with main effects and no interactions), and ELAAN-I. For example, the top panel in Figure 6 shows that among the tracts in the first quintile of the actual self-response rates, $85.1\%$ are correctly predicted by ELAAN-I to fall in the first quintile. The corresponding proportion for Lasso is $84.4\%$. The same panel shows that among the tracts in the first quintile of the actual self-response rates, $14.1\%$ are incorrectly predicted to fall in the second quintile; this proportion is smaller than the one for the Lasso ($14.7\%$). Similarly, in the second quintile of the actual self-response rates, a higher proportion ($58.4\%$) is correctly identified by ELAAN-I to be in the second quintile; this is again an improvement over the $56.6\%$ for the Lasso regression model. For all 5 quintiles of the actual self-response rates, ELAAN-I identifies a higher proportion to be in the correct quintile than Lasso, suggesting that our approach has an edge over the Lasso.

### v5

@@ 3 Incorporating strong hierarchy constraints
This section discusses the model with sparse interactions under strong hierarchy.
@@
For example, based on our analysis, covariates “Males” and “Females” appear to be unnecessary for self-response prediction when the strong hierarchy constraint is imposed via model (11), while a related covariate “Households with female without spouse” remains important. In contrast, the LRS uses both “Males” and “Households with females without spouse” (Erdman and Bates, 2016) – both features appear to be important perhaps due to the use of a linear regression model. Similarly, covariates pertaining to civilian employment status across different segments/age-groups of the population are excluded from our nonlinear AM when we learn the (sparse) main and interaction effects under a strong hierarchy.
@@
To our knowledge, we present a novel computational model to study estimator (12). The methodology presented here is of independent interest in the context of structured nonparametric learning with interactions. We refer to this model as ELAAN-H [^13] ELAAN-H stands for End-to-end Learning Approach for Additive spliNes with Hierarchy.. In the Supplement (Ibrahim et al., 2024), we propose algorithms to obtain good solutions to Problem (12).
@@
As mentioned earlier, the EBM approach (Nori et al., 2019) does not support hierarchy constraints. GAMI-Net (Yang, Zhang and Sudjianto, 2021) follows a two-stage approach to fit nonparametric additive models with sparse (weak) hierarchical interactions. GAMI-Net performs a screening step after estimating the main effects with neural-network blocks and only consider interaction effects that satisfy the weak hierarchy principle amongst the screened main effect – an interaction effect can appear in the model if one of the main effects is in the model. As mentioned earlier, GAMI-Net is not based on a sparsity-inducing penalized optimization procedure; and can be computationally much more expensive than our estimators.
@@ 4.1 Sparse additive model with interactions
We measure prediction accuracy using the integrated squared error, $\text{ISE}=\mathbb{E}_{\mathbf{x}}[(\hat{f}(\mathbf{x})-f^{*}(\mathbf{x}))^{2}]$, estimated by Monte Carlo integration using $10,000$ test observations from the same distribution as the training ones, following the procedure in Li, Lue and Chen (2000). We vary the number of training observations from $100$ to $400$ and use $100$ replications for each simulation setting. We compare our estimators to well-known benchmarks EBM (Nori et al., 2019), GAMI-Net (Yang, Zhang and Sudjianto, 2021), COSSO (Lin and Zhang, 2006), and MARS [^14] MARS is a stepwise forward–backward procedure for building functional ANOVA models. (Friedman, 1991).
Table 2 presents the average test ISE and their standard errors (based on the replications). The tuning parameters for ELAAN-I, ELAAN-H, EBM and GAMI-Net are selected via 5-fold cross-validation. The results for MARS and COSSO are taken from Table 6 in Li, Lue and Chen (2000). Table 2 demonstrates that both of our estimators, ELAAN-I and ELAAN-H, appear to work well compared to competitors. Due to space constraints, additional simulations comparing our approach to EBM and with similar conclusions are provided in the Supplement (Ibrahim et al., 2024).
@@
Next, we study the support recovery performances of ELAAN-I, ELAAN-H, EBM and GAMI-Net. We evaluate models using the discrepancy between the true and estimated support and separately report the average F1-score (defined in the Supplement (Ibrahim et al., 2024)) for the main and the interaction effects. The support recovery metrics are shown in Table 2. We observe that, overall, both ELAAN-I and ELAAN-H appear to work better compared to EBM and GAMI-Net in F1-score. Moreover, ELAAN-H is seen to be the best-performing method across all the simulation settings. In summary, our approaches work well in terms of variable selection.
@@
Table 3 presents additional variable selection details of the two best-performing approaches: ELAAN-I and ELAAN-H. For each true main and interaction effect, we report the relative frequency of their appearance in the estimated support across all replications. The results in Tables 2 and 3 suggest that ELAAN-H has an edge over ELAAN-I in terms of the variable selection when the true model is hierarchical. In practice, when we do not know if the underlying truth obeys a strong hierarchy, ELAAN-H tends to result in more compact models, as we demonstrate in the Census application in Section 5.
@@ 4.2 Large-scale setting with correlated features
ELAAN-H ${}~{}10.6\pm 1.8$ ${}~{}97.57\pm 0.68$ $97.57\pm 0.68$ $91.31\pm 1.52$
We produce 10,000 training observations and evaluate the prediction performance as before, using ISE on a test set of size 10,000. We compare our estimators to EBM (Nori et al., 2019) and GAMI-Net (Yang, Zhang and Sudjianto, 2021). Table 4 presents the average prediction errors and support recovery metrics over 25 simulation replications, along with the corresponding standard errors. The tuning procedure for ELAAN-I, ELAAN-H, EBM and GAMI-Net is described in the Supplement (Ibrahim et al., 2024). Table 4 demonstrates that our proposed models exhibit strong performance compared to competitors, showing a 7-11 fold reduction in the prediction error.
@@ 5 Case study: Predicting the Census Survey Self-Response Rate
We now present our findings on the Census application. As mentioned in Section 1, obtaining interpretable models with good predictive capabilities is important in this application – such models can help inform the planning of outreach campaigns and guide stakeholders in deciding the allocation of spending for different communications channels (for example, TV, radio, digital), advertising messages with appropriately tailored content, and the timing of spending during the campaigns (Kulzick et al., 2019). Due to their opaque nature, predictive models that topped the nationwide competition were not actionable for this application. On the other hand, the linear models that drive the LRS have limited predictive power. We explore how our proposed approach balances simplicity and good prediction performance.
The data we use for this study is publicly available in the US Census Planning Database, which provides a range of demographic, socioeconomic, housing, and Census operational data (US Census Bureau, 2019). The data in the Planning Database includes covariates from the 2010 Census and the 2013-2017 American Community Survey (ACS), aggregated at both the Census tract level and the block level. We use tract-level data, with approximately $74,000$ observations and $500$ covariates. The response is the ACS self-response rate. We exclude the following covariates from our model: spatial covariates (“State”, “County”, “Tract”, “Flag”, “AIAN Land”) and variables that serve as a proxy to the response (for example, “Low response score”, “Number of housing units that returned first forms”, “Replacement forms” or “Bilingual forms” in Census 2010). We also remove the margin of error variables corresponding to the ACS. After excluding these variables, we are left with $p=295$ covariates [^15] To clarify, the results in this section are based on these $295$ features, though our algorithms scale to $p=500$..
@@ 5.1 Experimental setup
Benchmark Methods. We evaluate and compare our models against existing linear and nonparametric approaches. We consider the following linear models with main effects only (i.e., without interactions): (i) Ridge regression; (ii) Lasso regression (Hastie, Tibshirani and Wainwright, 2015); (iii) $\ell_{0}$-penalized regression with additional ridge regularization (denoted as $\ell_{0}-\ell_{2}$), implemented via L0Learn (Hazimeh and Mazumder, 2020a). In addition, we consider the following linear models with interaction effects: (iv) Lasso with main effects and all pairwise linear interactions, (v) hierScale: a convex optimization framework for learning sparse main effects and interactions under a strong hierarchy constraint (Hazimeh and Mazumder, 2020b). For (i), (ii), and (iv), we use Python’s scikit-learn library (Pedregosa et al., 2011).
In addition to the competing methods discussed above based on linear models, we consider nonparametric additive models with/without interactions. For AMs with main effects only, we compare with (vi) Additive Models with Lasso-based selection, which we fit using Python’s scikit-learn library (Pedregosa et al., 2011). For additive models with interactions, we consider (vii) EBM: A boosting approach that uses trees for main and interaction effects (Nori et al., 2019). (viii) GAMI-Net: a neural network-based additive model with structured interactions (Yang, Zhang and Sudjianto, 2021). For EBM, we tune the learning rate in the range $[10^{-4},10^{-1}]$ and the number of interactions in $[10,500]$ for 1000 trials. For GAMI-Net, we use a $200$ batch-size, $0.001$ learning rate, $500$ interaction effects, $500$ epochs (for each stage) and $0.001$ loss threshold; we set the NN basis to 2-layered network with 10-neurons. [^16] The default NN basis of 5-layers with 40 neurons per-layer had long runtimes (3-days on a CPU), so a referee suggested the alternative hyperparameters. See the code in supplementary files (Ibrahim et al., 2024) for more details.
@@
All our algorithms for estimators (a)–(c) are implemented in Python. We use cubic B-splines with $10$ knots for the main effects. For the interaction effects, we use tensor spline bases of degree $3$ with 5 knots in each coordinate, leading to a total of $5\times 5=25$ knots [^17] We study the effect of number of knots on out-of-sample generalization on synthetic data (see the Supplement (Ibrahim et al., 2024)).. Because the problem at hand has $295$ covariates and $43,365$ possible pairwise interactions, we need to be careful with implementation aspects while generating spline-transformed representations for all the interaction effects, which can be memory intensive.
For ELAAN and ELAAN-I we perform a tuning procedure with warm-starts over a 2D grid of parameters ($\lambda_{1}$, $\lambda_{2}$) – for details, see Section S1.4.3 of the Supplement (Ibrahim et al., 2024). For both ELAAN-I and ELAAN-H, we set $\alpha=1$ so that the main and the interaction effects have the same $\ell_{0}$-penalty parameter. For ELAAN-H, we fix the smoothing parameter $\lambda_{1}$ to the optimal value available from ELAAN-I. Parameter $\lambda_{2}$ (as well as parameter $\tau$ defined in Section S3 of the Supplement (Ibrahim et al., 2024)) is chosen based on validation tuning. For all our methods (ELAAN, ELAAN-I, ELAAN-H), we cap the number of tuning-parameter values at 1000.
@@ 5.2 Comparing methods: prediction, sparsity and model structure
Table 5: Comparisons of our methods with several benchmark models as discussed in the text. We display the average test RMSE for the different models, along with the corresponding number of covariates, and a number of effects (main and interactions). The best metrics are highlighted in bold. Numbers after $\pm$ provide the standard errors. Asterisk(*) indicates statistical significance (p-value¡0.05) over the best existing model, using a one-sided paired t-test. Models (ix), (x) are black-box models using higher order interactions, hence #Main and #Interactions are left as ‘-’ (dash).
@@
Additive models vs black-box ML methods. Table 5 reports the prediction errors (RMSE) on the test set along with the number of nonzero features in the model. Importantly, we observe that the test performances of AMs with sparse interactions (both with and without the hierarchy constraints) are better than that of the best black box predictive ML models (ix), (x) that are based on GBDT, Neural Networks. Our model ELAAN-H delivers the best RMSE value and is closely followed by ELAAN-I.
ELAAN-I/ELAAN-H  vs state-of-the-art for AMs with sparse interactions. Among the methods we compared, EBM and GAMI-Net were the only methods that could compute AMs with sparse interactions for the Census dataset. Interestingly, the test performance of both ELAAN-I and ELAAN-H is better than that of EBM and GAMI-Net. Interpretability (i.e., model sparsity and/or hierarchy) is another key factor differentiating the leading prediction methods. Our models select a smaller number of additive components than EBM and GAMI-Net. For example, ELAAN-I selects a total of 234 additive components. In contrast, EBM and GAMI-Net select 787 and 325 additive components, respectively. We also observe that our models are almost twice as compact in terms of the number of selected covariates.
@@
We briefly discuss computation times for different methods. EBM takes around one hour to fit a model with 500 pairwise interactions (for a single tuning parameter). Computing a GAMI-Net solution (for one tuning parameter corresponding to 500 interaction effects with a 2-layered NN basis per effect) takes around $12$ hours [^18] The default implementation with a 5-layered NN basis per effect took 3 days on a CPU. using 8 CPUs and 3 hours using a single V100 Tesla GPU. In contrast, our approaches are much faster: ELAAN-I takes on average $1.8$ mins to compute a solution for a fixed tuning parameter [^19] We use warm-start continuation across $\lambda_{2}$-values for a fixed $\lambda_{1}$: it takes approximately 3 hours to compute 100 solutions where the maximum number of pairwise interactions is set to 500.. When ELAAN-H is run on a reduced subset of 800800~{}800800 pairwise interactions, it takes $1.5$ hours to compute a path of 100 solutions. These numbers are reported on an 8-CPU 128GB RAM device.
@@
Nonlinear models vs linear models. Table 5 suggests that nonlinear/nonparametric models have better predictive performance than their linear counterparts. Also, linear models with sparse main and interaction effects appear to have an edge over sparse linear models without interactions. Another appealing aspect of sparse nonparametric AMs compared to linear models, is in model parsimony – we alluded to this aspect in Figure 3 while using a reduced number of features. The number of nonzero effects is significantly lower for the nonlinear AMs. For example, $\ell_{0}$-sparse nonparametric AMs with no interactions (i.e., ELAAN) can achieve the same level of predictive performance as its linear model counterpart (i.e., Model (iii)) with significantly fewer covariates: $40$ versus $145$. Similarly, if we compare hierScale (i.e., linear main and interaction effects with strong hierarchical restrictions) to its nonparametric counterpart i.e., ELAAN-H, the number of main effects (and also, covariates) reduces from $276$ to $133$, and the number of interaction effects reduces from $4,107$ to $255$. Interestingly, all the covariates selected by the nonparametric AMs with interaction models (with/without strong hierarchy) are contained in the set of covariates selected by the sparse linear models with both main and interaction effects.
@@
Additive nonlinear models: interactions vs no interactions. Table 5 shows that ELAAN (AMs with no interactions) has more covariates than AMs with interactions, both ELAAN-I and ELAAN-H. By including interactions that obey the hierarchy principle, we select a smaller number of covariates: ELAAN-I has 154 covariates vs ELAAN-H has 133. The difference between ELAAN vs ELAAN-I is possibly because many nonlinear main effects attempt to explain the nonlinear interaction effects. This reduction points to the redundancy of some of the covariates when interaction effects are directly included in the model. The prediction performance of nonlinear AMs seems to improve quite a lot when pairwise interactions are included in the model.
@@
Hierarchy vs no hierarchy. We observe that ELAAN-H, i.e., the additive model with hierarchical interactions, achieves the best out-of-sample RMSE – this improves over ELAAN-I (interactions with no hierarchy). The improvement is statistically significant based on a paired t-test. We observe that ELAAN-I selects fewer effects (33 Main + 201 Interactions) when compared to ELAAN-H (133 Main + 255 Interactions). On the other hand, ELAAN-H obeys the hierarchy constraint and selects fewer covariates than ELAAN-I– this can aid in interpretability and be easier to operationalize from a practical standpoint.
@@
Insights from visualizations. As illustrated by Figure 2, which displays some marginal nonlinear fits, an appealing aspect of nonlinear AMs is that they naturally allow the practitioner to gather insights into the associations between the response and a feature of interest, for example, by exploring the map $x_{j}\mapsto f_{j}(x_{j})$. Similarly, the map $(x_{j},x_{k})\mapsto f_{j,k}(x_{j},x_{k})$ would shed light into how the interaction of $(x_{j},x_{k})$ influences the output. Such interpretations are a by-product of our additive model framework, and may not be readily available via black-box ML methods such as Neural Networks and GBDT—see also Lou et al. (2013) for related discussions advocating for the interpretability of AMs. These association plots, for example, can help stakeholders identify promising factors influencing self-response scores and potentially informing policy decisions (e.g., targeted investments and optimizing operational costs).
@@
To obtain a finer understanding of the performance of our models, we use visualization tools inspired by earlier works from the US Census Bureau (US Census Bureau, 2017; Kulzick et al., 2019). Figure 4 illustrates the tract self-response rates predicted by ELAAN-I i.e., model (8) on a map of the United States. Different from Figure 1, which shows actual data with missing values for some of the tracts [^20] Some responses are deliberately suppressed by the US Census Bureau for privacy considerations, to limit the disclosure of information about individual respondents and to reduce the number of estimates with unacceptable levels of statistical reliability (US Census Bureau, 2016), Figure 4 provides predicted self-response rates for all tracts based on our proposed model. Predictive models in general and our models in particular, allow us to forecast the response rates for tracts where gathered data is incomplete. It is worth noting that our regression-based models may be preferable over commonly used nearest neighbor type methods: the latter models may not readily offer insights into factors associated with high/low response rates. Nearest neighbor-based methods may also be ill-posed for isolated regions such as Alaska, Hawaii, and Puerto Rico, each having at least $10\%$ of their tracts censored.
@@
Figure 6 shows how the sorting of the Census tracts into quintiles (US Census Bureau, 2017; Kulzick et al., 2019) of the actual self-response rates compares with the corresponding sorting of the predictions made by: the Lasso (we use an $\ell_{1}$-penalized linear model with main effects and no interactions), and ELAAN-I. For example, the top panel in Figure 6 shows that among the tracts in the first quintile of the actual self-response rates, $85.1\%$ are correctly predicted by ELAAN-I to fall in the first quintile. The corresponding proportion for Lasso is $84.4\%$. The same panel shows that among the tracts in the first quintile of the actual self-response rates, $14.1\%$ are incorrectly predicted to fall in the second quintile; this proportion is smaller than the one for the Lasso ($14.7\%$). Similarly, in the second quintile of the actual self-response rates, a higher proportion ($58.4\%$) is correctly identified by ELAAN-I to be in the second quintile; this is an improvement over the $56.6\%$ for the Lasso regression model. For all 5 quintiles of the actual self-response rates, ELAAN-I identifies a higher proportion to be in the correct quintile than Lasso, suggesting that our approach has an edge over the Lasso.